Find the position of center of mass of a system of 3 particles of masses 1kg, 2kg, 3kg that are located at 3 corners of an equilateral triangle of side 1m. Take 1kg mass at the origin and 2kg along x-axis.

Grade:11

1 Answers

Arun
6 years ago

To find the center of mass:

  1. Fix one vertex of triangle as origin and give coordinates to other 2 vertices.
  2. apply the center of mass formula

A(0, 0) , B (1,0) and C(1/2, sqrt(3)/2)

Xcm = M1X1+M2X2+M3X3/(M1+M2+M3) = 1(0)+2(1)+3(1/2)/(1+2+3) = 7/12

Ycm = M1Y1+M2Y2+M3Y3/(M1+M2+M3) = 1(0)+2(0)+3(sqrt(3)/2)/(1+2+3) = sqrt(3)/4